OpenStudy (anonymous):

write the rate in lowest terms .... Jan earned $62 in 12 hours

OpenStudy (anonymous):

Lowest terms means to put the problem into $ x / 1 hr correct?

OpenStudy (anonymous):

figure the rate and then put it in lowest terms

OpenStudy (anonymous):

192/6 = 10/x

OpenStudy (anonymous):

In this case you would need to multiply by x on both sides of the equation to get the x to be on the top. Then can you solve it?

OpenStudy (anonymous):

So you would have x * 192 / 6 = 10

OpenStudy (anonymous):

Arfney, can you just please give me the answer? I have to hurry. Please